A skyrocket is shot into the air. It's altitude h in feet/seconds is given by the function h = -16t^2 + 128t.
a. In how many seconds does the skyrocket reach its maximum altitude?
b. What is the skyrocket's maximum height?

Answer by CubeyThePenguin(3113)

A) max height at -b/2a = 128/32 = 4 seconds

B) plug in t = 4

-16(4)^2 + 128(4) = -16(16) + 512 = 256 feet